Here is the problem: Dordine Company acquired equipment on March 1, 2020, at a cash cost of $1,000,000. Transportation charges amounted to $12,000, installation cost was $40,000, testing costs were $30,000, and the payment of a fine for not getting the proper permit for moving the equipment was $1,000. The equipment was estimated to have a useful life of 12 years and a salvage value of $60,000 at the end of its life. It was further estimated that the equipment would be used in the production of $364,000 units of product during its life. During the 2020 year, 40,000 units of product were produced. Required: A. Prepare a schedule to calculate the depreciable base of the asset. B. Compute the depreciation to the nearest dollar for the year ended December 31, 2020, using the straight-line method. C. Compute the depreciation to the nearest dollar for the year ended December 31, 2020, using the units-of-production method. D. Compute the depreciation to the nearest dollar for the year ended December 31, 2020, using the sum-of-the-years-digits method. E. Compute the depreciation to the nearest dollar for the year ended December 31, 2020, using the double-declining balance method.
Answer #1

Question A

The cost of a fixed asset comprises of its purchase price, including import duties and other non-refundable taxes or levies and any directly attributable cost of bringing the asset to its working condition for its intended use.

Examples of directly attributable costs are :-

(i) site preparation

(ii) initial delivery and handling costs

(iii) installation cost, such as special foundations for plant

(iv) professional fees, for example fees of architects and engineers.

Therefore the Cost of the Equipment for the company will be as follows:-

Cash Cost = $1,000,000

Transportation charges = $12,000

Installation cost = $40,000

Testing cost = $30,000

Payment of Fine for moving the equipment ($1,000) will not be considered as the item to be capitalized for arriving at the value of the equipment as the same is a fine or penalty which is born because the permit was not proper therefore the same will be attributed to expense.

Total cost = $1,082,000

Depreciation base to the asset will be (Total cost) = $,1,082,000 - $60,000 = $1,022,000

Question B.

Computation of depreciation as per Straight line method for the year ended December 31, 2020

Cost of the equipment $10,82,000.00
Salvage Value $60,000.00
Useful life of the asset (Years) 12
Depreciation per year for Dec 31, 2020   $85,166.67
Calculation =($10,82,000-$60,000)/12

Question C.

Computation of Depreciation as per Units-of-production method for the year ended December 31, 2020

= [(Original Value − Salvage Value)/Estimated Production Capability​] × Units per year

Cost of the equipment $10,82,000.00
Salvage Value $60,000.00
Expected number of Units to be produced in year 2020   3,64,000
Depreciation per year for Year ended Dec 31, 2020 $1,12,307.69
Calculation ={($10,82,000-$60,000)/3,64,000}*40000  

Question D.

Computation of depreciation as per Sum of years digit method for the year ended December 31, 2020

Cost of the equipment $10,82,000.00
Salvage Value $60,000.00
Expected Useful Life (Years) 12
Sum of Years =1+2+3+4+5+6+7+8+9+10+11+12 = 78
Remaining Years in 2020 12
Percentage of depreciation in year 2020 =12/78 = 15.38%
Depreciation Calculation =($10,82,000-$60,000) * 15.38%
Depreciation for the year ended Dec 31, 2020 $1,57,183.60

Question E.

Computation of Depreciation as per the Double declining balance method for the year ended December 31, 2020

Formula for depreciation rate = Straight line method rate * 2 * Book Value

Straight line Rate = $85,166.67/$10,82,000 = 7.87%

Depreciation rate as per double declining balance method = 7.87% * 2 = 16.74%

Depreciation as per double declining balance method for the year ended Dec 31, 2020 = $10,82,000 * 16.74% = $1,70,333.34